OpenStudy (djbreezy):

Which of the ordered pairs in the form (x,y) is a solution of this equation? 5x-y/3=13 (2,-9), (3,-6) A. Both are solutions B. Neither is a solution C. The first is a solution,but the second is not. D. The first is not a solution but the second is.

OpenStudy (djbreezy):

@isaidavila

OpenStudy (isaidavila):

Replace First (2,-9) 5(2)-(-9)/3=13 10+3=13 13=13 is correct so is a solution Now (3,-6) 5(3)-(-6)/3=13 15+2=13 17=/13 so in not a solution

OpenStudy (djbreezy):

So the answer is C @isaidavila